OpenStudy (anonymous):

Aki’s bicycle has determined that when x hundred are built, the average cost per bicycle is given by c(x) = 0.5x^2- 1.9x+10.45, where c(x) is an hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bicycle?

OpenStudy (anonymous):

derive it, find critical points, determine if derivative is going from negative to positive at each point. If so, voila, minimum

OpenStudy (anonymous):

I do not get what you are saying could you please explain

OpenStudy (anonymous):

sorry is this for algebra 2 or calc class?

OpenStudy (anonymous):

Algebra

OpenStudy (anonymous):

if its for algebra 2, do it the lame way and do x = -b/2a at the vertex

OpenStudy (anonymous):

Thank you any way I still do not know what you are tell me I guess I will try back tomorrow

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

a = .5, b = -1.9, plug those into -b/2a, and you'll get the answer

OpenStudy (anonymous):

I am sorry I have been out of school for a while and am earning a college degree and this is all new to me again thank you but I will try again tomorrow

OpenStudy (anonymous):

the answer is x = 190 bicycles.

OpenStudy (anonymous):

Ok so how did you get that answer did you divide the -1.9 by 2(5)

OpenStudy (anonymous):

so you would divide the -1.9 divided by 10

OpenStudy (anonymous):

2(.5) =1 so -1.9/1 = -1.9 but it is -b/2a, so multiply that times -1 to get x= 1.9

OpenStudy (anonymous):

a = .5

OpenStudy (anonymous):

and each x is 100 bicycles, so 1.9 *100 = 190 bikes

OpenStudy (anonymous):

Oh Ok thank you now I get